OpenStudy (anonymous):

What is the equation of a parabola with a vertex at (0, 0) and a focus at (0, 3).

OpenStudy (anonymous):

Do you have any idea of how to come across this?

OpenStudy (anonymous):

no

OpenStudy (psymon):

For a parabola, we either have the equation x^2 = 4py or y^2 = 4px. The focus will always be inside of the parabola, too. So if the vertex is (0, 0) and the focus is at (0,3), the parabola must be opening upward. When the parabola opens upward, we have the equation x^2 = 4py. When the equation is of the form x^2 = 4py, the focus is at the point (0, p) That being said, p must = 3 since the focus is (0,3). So plugging in 3 for p, the equation is x^2 = 12y

OpenStudy (anonymous):

Well thanks for saving me the written explanation lol @Psymon

OpenStudy (anonymous):